Air is pumped into a spherical balloon such that the surface area of the balloon increases at a rate of 10 cm2 per second. Find the rate at which the radius is increasing when the radius is 4 cm. [ i.e. find when r = 4] [Note : surface area, A = 4πr2 ]


1
Expert's answer
2021-09-09T09:17:01-0400

We know that , surface area of spherical balloon "(A)=4\pi r^2"

Differentiating both sides w.r.t "t" , we get

"\frac{dA}{dt}=\frac{d}{dt}(4\pi r^2)
\\\frac{dA}{dt}=4\pi\frac{d}{dt}( r^2)
\\\frac{dA}{dt}=8\pi r\frac{dr}{dt}...(1)"

Now, it is given that

"\frac{dA}{dt}=10\ cm^2/s" and we need to find "\frac{dr}{dt}" at "r=4" .

So, from equation (1), we get

"10=8\pi \times 4\times \frac{dr}{dt}
\\10=8\times\frac{22}{7} \times 4\times \frac{dr}{dt}
\\\frac{dr}{dt}=\frac{35}{352}=0.994 \ cm/s"